OpenStudy (anonymous):

You have a CD with 14 cuts on it, and another with 11 cuts by a different artist. You wish to make a compilation from all these songs containing just 3 different cuts which you will play for your friends. How many possible ways can you arrange 3 cuts copied from these two CDs? (Count two compilations as different if they have the same three songs in a different order.) A. 25 B. 13,800 C. 2,300 D. 17,550 E. 180

OpenStudy (kropot72):

There is a total of 25 cuts. The number of permutations of these cuts taken 3 at a time is:\[\frac{25!}{(25-3)!}=25\times 24\times 23\]

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

:)